# A Survey of Deep Learning for Geometry Problem Solving

### v2

@@ 5.1 Challenges
Capability. Current deep learning systems still have many deficiencies in their capabilities when it comes to solving geometry problems. Given that most geometry problems involve multimodal input, the model’s geometric Visual Perception ability is particularly important. However, studies have shown that after introducing geometric figures, the model’s accuracy is lower than when only the geometry problem text is input Zhang et al. (2025e); Onuoha et al. (2025). For multimodal geometry problems, the model’s spatial perception of geometric diagrams has always been a key bottleneck, seriously restricting the improvement of problem-solving performance Sun et al. (2024); Xing et al. (2024); Kamoi et al. (2024); Zhang et al. (2024c). Some studies have pointed out that deep learning models are deficient in the recognition of geometric angles Okada et al. (2023); Cho et al. (2025a) and perception Wang et al. (2025d); Weng et al. (2025). Other studies believe that the model is still lacking in the perception of line length Wei et al. (2024); Huang et al. (2025b).This may be because although models have the ability to solve geometry problems, their capabilities are limited by a one-dimensional sequential architecture, and the spatial perception that humans can intuitively understand is difficult for models to grasp Sun et al. (2025d). It may also be because the image resolution of the visual encoder is still low Zhang et al. (2024a); Zhu et al. (2025) and it is mainly trained based on natural images Hsu et al. (2022); Sharma et al. (2025), resulting in its suboptimal performance when processing geometric figures. In addition, current deep learning models still have deficiencies in Arithmetic Accuracy. Some models use symbolic or formal analysis methods to deal with problems Ning et al. (2025); Chervonyi et al. (2025), while others introduce external computing modules to run the code to alleviate this problem Duan et al. (2024); Zhang et al. (2024f); Pan et al. (2024). LLMs may also develop a Mindset when solving geometry problems, such as their tendency to build coordinate systems first Sun et al. (2025d). However, it may fail when this general approach no longer applies.
@@ 6 Conclusion
In this paper, we provide a comprehensive and systematic survey of geometry problem solving. We summarize the relevant tasks, deep learning architectures and other methods, evaluation metrics and methods in detail, and provide an in-depth analysis of the shortcomings of existing data, evaluation, and model capabilities. Finally, we look forward to possible future research directions in this field and emphasize the vast space for exploration in this field. This article aims to provide readers who are interested in this field with a comprehensive and practical resource to meet their research needs.

### v3

@@ 5.1 Challenges
Capability. Current deep learning systems still have many deficiencies in their capabilities when it comes to solving geometry problems. Given that most geometry problems involve multimodal inputs, models’ geometric visual perception capabilities are particularly crucial. However, studies have shown that introducing geometric figures often decreases model accuracy compared to using text input alone Zhang et al. (2025e); Onuoha et al. (2025). In multimodal geometry problems, spatial perception of geometric diagrams remains a major bottleneck, severely limiting problem-solving performance Sun et al. (2024); Xing et al. (2024); Kamoi et al. (2024); Zhang et al. (2024c). Some studies point out that deep learning models struggle with detecting Okada et al. (2023); Cho et al. (2025a) and perceiving Wang et al. (2025d); Weng et al. (2025) geometric angles. Others suggest that models also lack accurate perception of line lengths Wei et al. (2024); Huang et al. (2025b). This may be due to the inherently one-dimensional, sequential nature of current model architectures, which limits their ability to grasp the kind of spatial perception humans intuitively possess Sun et al. (2025d). Another reason could be the low resolution of visual encoders Zhang et al. (2024a); Zhu et al. (2025), which are primarily trained on natural images Hsu et al. (2022); Sharma et al. (2025), leading to suboptimal performance on geometric figures. Furthermore, many deep learning models also struggle with arithmetic accuracy. Some models adopt symbolic or formal reasoning methods Ning et al. (2025); Chervonyi et al. (2025), while others incorporate external computation modules to execute code and mitigate this issue Duan et al. (2024); Zhang et al. (2024f); Pan et al. (2024). In addition, LLMs may also exhibit a certain mindset when solving geometry problems, such as a tendency to construct coordinate systems as a default strategy Sun et al. (2025d). However, this approach can fail when it does not fit the context of the specific problem.
@@ 6 Conclusion
In this paper, we provide a comprehensive and systematic survey of geometry problem solving. We summarize the relevant tasks, deep learning architectures and other methods, evaluation metrics and methods in detail, and provide an in-depth analysis of the shortcomings of existing data, evaluation, and model capabilities. Finally, we look forward to possible future research directions and emphasize the vast space for exploration in this field. This article aims to provide readers who are interested in this field with a comprehensive and practical resource to meet their research needs.